MODULE IV: TIME VALUE OF MONEY INTRODUCTION The time value of money analysis has many a lysis has many applications, rangingPresent Value Formula: Excel: PV = FVN/(1+1)N =PV(I, N, PMT, FV, type) Example 3: What is the present value of $500 to be recWhat is the present value of a 3-vear annuity with vear given as percent annual interest rate? lue of a 3-year annuity with pfollow. Assuming an annual discount rate of 10 percent, what is the ne this investment? le net present value of Answer: If thTime Value of Money Exercise: Question 1: Assume you deposit $700 every three months at ercent annual rate, compounded $700 eTime Value of Money Spreadsheet Example ASSIGNMENT # NAME: LAB TIME: Answer Sheet for Time Value of Money Exercise From Previ

I need help on question 2.
Answer #1

Question 2]

See Example 8

The monthly rate is calculated using RATE function in Excel :

nper = 60 (5 year loan with 12 monthly payments each year. Total number of payments = 5 * 12 = 60).

pmt = 250 (monthly payment)

pv = -13000 (Amount borrowed)

RATE is calculated to be 0.4817%

А14 ro f =RATE(60,250,-13000) J K L А B C 14 |0.4817%

This is the monthly rate. To get APR, we multiply this by 12.

APR = 5.78%.

А14 f =RATE(60,250,-13000)*12 J K L B C 14 | 5.78%|